# EDTformer: An Efficient Decoder Transformer for Visual Place Recognition

### v1

@@ III-C Low-rank Parallel Adaptation
The foundation model DINOv2 recently has become a popular backbone for feature extraction in VPR, generally accompanied by a fine-tuning/adaptation for better performance. CricaVPR [24] and SelaVPR [25] have used the PETL methods (i.e., inserting tunable adapters into the frozen encoder blocks) to adapt DINOv2, as shown in Fig. 3 (b). However, they are efficient only in terms of parameters, not in training time and memory usage. Here, we conduct a simple theoretical derivation to explain this point. For a network with $L$ blocks, each having output ${z}_{i}$, a large number of frozen parameters $\theta_{i}$ and a few trainable parameters $\psi_{i}$, our goal is to minimize a loss function, $\mathcal{L}$ based on stochastic gradient descent. Concretely, to update $\psi_{i}$ during backpropagation, we first need to compute $\frac{\partial\mathcal{L}}{\partial\psi_{i}}$ via the chain rule as follows
$\frac{\partial\mathcal{L}}{\partial\psi_{i}}=\underbrace{\frac{\partial\mathcal{L}}{\partial z_{L}}\frac{\partial z_{L}}{\partial z_{L-1}}\cdots\frac{\partial z_{i+1}}{\partial z_{i}}}_{\text{intermediate activations}}\frac{\partial z_{i}}{\partial\psi_{i}}=\frac{\partial\mathcal{L}}{\partial z_{i}}\frac{\partial z_{i}}{\partial\psi_{i}}.$ (8)
Although $\frac{\partial z_{i}}{\partial\psi_{i}}$ can be directly computed, in order to calculate $\frac{\partial\mathcal{L}}{\partial z_{i}}$, it is inevitable to compute gradients about the intermediate activations, as shown in Eq. (8), which can incur huge memory overhead and additional running time.
To tackle the above issue, we propose the Low-rank Parallel Adaptation method inspired by previous studies [51, 52]. Instead of inserting tunable parameters inside the backbone, we design a parallel network that does not require backpropagation gradients through the frozen backbone during training as shown in Fig. 3 (c). It is a lightweight and separate network, which directly takes the intermediate features from the backbone as inputs and refines them progressively to produce more accurate representations.
@@ III-D Training Strategy
The proposed model is trained on the GSV-Cities [36] dataset following its standard framework. GSV-Cities contains 560k images with highly accurate labels depicting 67k different places. We apply the multi-similarity (MS) loss [37] function with an online hard mining strategy. The MS loss is formulated as
@@ IV-A Datasets and Evaluation Metrics
To demonstrate the effectiveness of our method, we conduct the experiments on multiple benchmark datasets, which present a diverse set of challenges encountered in the real world. Table I provides a concise summary of these benchmark datasets. Pitts30k [38], extracted from Google Street View, shows significant changes in viewpoint. MSLS [39] is collected from 30 major cities across six continents over nine years period, covering all seasons and encompassing diverse weather, cameras, daylight conditions, and structural settings. Tokyo24/7 [40] presents severe lighting (day/night) variations. SPED [60] is collected from surveillance cameras and encompasses a diverse array of natural scenes. Nordland [41] is captured using a front-mounted train camera, spanning all four seasons. We use the winter images as queries and summer images as database followed by previous work [1, 45, 24]. AmsterTime [42] exhibits substantial image modality variations, using historical grayscale images as queries and contemporary color images as database. SVOX [43] is a cross-domain VPR dataset gathered under diverse weather and lighting conditions. We primarily utilize the three most challenging query subsets: SVOX Night, SVOX Rain and SVOX Sun.
@@ IV-B Implementation Details
[…]
@@ IV-C Comparison with State-of-the-art Methods
In this section, we compare our proposed method with a wide range of existing SOTA VPR algorithms, including eight single-stage VPR methods using global features for direct retrieval: NetVLAD [12], SFRS [15], CosPlace [44], MixVPR [33], EigenPlaces [45], CricaVPR [24] and BoQ [58], as well as three superior two-stage VPR methods: Patch-NetVLAD [16], $R^{2}$ormer [17] and SelaVPR [25]. Notably, our work uses the same training dataset as MixVPR, CricaVPR, SALAD and BoQ, i.e., GSV-Cities. Additionally, the latest works CricaVPR, SALAD, BoQ and SelaVPR all leverage the foundation model DINOv2 as the backbone (SelaVPR using DINOv2-large, while others using DINOv2-base) and achieve the SOTA performance on multiple benchmarks. Thus, we follow them and apply the DINOv2-base. Table II shows the quantitative results on MSLS, Tokyo24/7, SPED and Pitts30k. Our method achieves the best R@1/R@5/R@10 performance on almost all datasets.
@@
Results analysis. CricaVPR, SALAD, BoQ, SelaVPR and our method all achieve superior performance on these datasets. Especially on Tokyo24/7, which shows severe illumination changes, CricaVPR, SALAD, BoQ and SelaVPR achieve 93.0% R@1, 94.6% R@1, 95.2% R@1 and 94.0% R@1 respectively. However, our method consistently improves performance on Tokyo24/7, achieving an incredible 97.1% R@1. This improvement stems from the robust and discriminative global descriptors generated by our model. The MSLS dataset is highly challenging, covering different seasons, weather and illumination conditions, various camera types and viewpoints, as well as different levels of dynamic objects presented in the scenes. Nevertheless, our method achieves 78.4% R@1, 89.8% R@5 and 91.9% R@10 on MSLS-challenge, showing significant advantages and outperforming other global-retrieval-based methods and two-stage methods by a considerable margin. Although our method achieves 92.0% R@1 on MSLS-val, 0.2% lower than SALAD, the dimensionality of our global descriptors is less than half that of SALAD. This indicates that our global descriptors can require less storage space and achieve a higher retrieval efficiency. In addition, we also get an overall better performance on the SPED dataset, demonstrating the high robustness of our method to handle various place images concerning natural scenes. On the Pitts30k dataset, which shows severe viewpoint changes, EDTformer outperforms all other methods but is slightly behind CricaVPR. This is because CricaVPR uses the cross-image encoder to combine the information of multiple query images from the same place in a batch at once in inference. As the number of query images in the batch decreases, especially when it becomes 1, the performance of CricaVPR drops significantly, as shown in Table II CricaVPR-1.
@@
To further evaluate the generalization of our method in some extreme scenarios, we conduct extensive experiments on other challenging datasets: Nordland, AmsterTime, SVOX Night, SVOX Rain and SVOX Sun. The results shown in Table III demonstrate the powerful capability of our method to effectively tackle the VPR task in difficult scenes. On the Nordland dataset, which exhibits significant variations in seasons and illumination, our EDTformer gets the best performance, even outperforming the two-stage SelaVPR. Additionally, our method surpasses other methods on the AmsterTime dataset by a large margin. Specifically, it outperforms the second by 11.6%, 10.8% and 8.8% on R@1/R@5/R@10 respectively. This implies the high robustness of our method to handle image modality variations in the datasets both containing grayscale and colorful images. On the SVOX Night, SVOX Rain and SVOX Sun datasets, our EDTformer achieves a relatively perfect performance (e.g., 98.7% R@1, 99.8% R@5 and 99.8% R@10 on SVOX Rain), which shows that it can be well applied to these special scenarios. In summary, the results on these challenging datasets further highlight the robust generalization capability of our model.
Moreover, our method also exhibits an advantage in terms of efficiency. Fig. 4 exhibits the R@1 performance on the Pitts30k dataset and the inference time of a single image about six global-retrieval-based methods, including SFRS, MixVPR, SALAD, BoQ, SelaVPR (global) and our EDTformer. MixVPR uses the CNN model (ResNet-50) as the backbone and proposes a feature mixing method to get global descriptors, which achieves the fastest inference speed. Although SFRS also leverages the CNN model (VGG-16), it applies PCA to reduce the dimensionality of the global descriptors, which is time-consuming. The other methods are all based on the vision foundation models. Specifically, SelaVPR (global) adopts DINOv2-large as the backbone, resulting in the slowest inference speed (16.95 ms). Among SALAD, BoQ and our EDTformer, all of which use DINOv2-base as the backbone, our method achieves the fastest inference speed (9.16 ms). In addition, we also provide the training memory consumption and training time per epoch for the three methods, as shown in Table. IV. Our training speed is approximately 15% faster and the training memory usage is much less than half of theirs. In summary, our method is efficient both in training and inference while achieving a superior recognition performance.
@@
Figure 4: The R@1 and inference time comparison of different single-stage methods on Pitts30k. We measure the inference time on an NVIDIA GeForce RTX 4090 GPU.
@@
TABLE IV: The comparison of SALAD, BoQ and our EDTformer in training memory usage and training time. We train the three models on GSV-Cities with batch size set to 36 and consistently measure the metrics on an NVIDIA GeForce RTX 4090 GPU.
Metric SALAD BoQ EDTformer
Training Memory Usage (GB) 14.81 15.21 5.72
Training Time (min / epoch) 11.21 10.67 8.77
@@ IV-D Ablation Studies
In this section, we first conduct a series of ablation experiments to demonstrate the effectiveness of our proposed EDTformer and LoPA. Subsequently, we further explore the impact of some design details about the EDTformer and LoPA. Notably, we consistently use the GSV-Cities dataset with the multi-similarity loss for training in the ablation experiments. Unless otherwise specified, the dimensionality of the global features produced by our method is 4096.
@@
Effect of EDTformer. To validate the effectiveness of the proposed EDTformer, we compare it with current common aggregation methods, including GeM [14], NetVLAD [12], Conv-AP [36], MixVPR [33] and BoQ [58] (related to our method). The “GeM-linear” represents the use of the GeM aggregator, followed by a linear layer to adjust the dimensionality of the descriptors. Likewise, “NetVLAD-PCA” implies that a PCA layer is integrated after the NetVLAD aggregator to perform dimensionality reduction. To ensure the fairness of the experiments, we consistently use DINOv2-base as the backbone with LoPA (rank set to 4) for feature extraction. The results are presented in Table V. When the dimensionality of global descriptors is the same (4096-dim), our EDTformer achieves highly competitive performance and outperforms some other aggregators (e.g., GeM, NetVLAD, Conv-AP and MixVPR) by a large margin. Although BoQ also achieves outstanding results, it is worth mentioning that the dimensionality of the global descriptors (12288-dim) produced by BoQ is more than three times of that of our global representations (4096-dim). In this case, our method still achieves a better performance than BoQ, especially on the Nordland dataset (obtaining an improvement of 4.5% in R@1). Moreover, we simultaneously present the recognition performance of our global descriptors with varying dimensions (achieved by only adjusting the last linear layer in our pipeline). As the dimensionality of the descriptors decreases, the overall performance gradually declines. However, the 512-dim descriptors can still achieve an impressive 90.1% R@1 on MSLS-val, 89.3% R@1 on SPED, 91.7% R@1 on Pitts30k and 71.1% R@1 on Nordland, surpassing the performance of the 768-dim descriptors outputed by GeM and remaining competitive against some SOTA methods such as SFRS [15], CosPlace [44] and EigenPlace [45]. Therefore, our EDTformer is also well-suited for some VPR scenarios that urgently require low-dimensional descriptors.
TABLE VI: Ablation on different fine-tuning/adaptation methods. “PartialTune-2” and “PartialTune-4” respectively denote only training the last 2 and 4 encoder blocks of DINOv2. “MCAdapter” is the abbreviation of MultiConvAdapter. The trainable parameters introduced by different methods are reported. We use an NVIDIA A800 GPU to measure the training memory usage with batch size set to 72.
Method Params $\downarrow$ Memory $\downarrow$ MSLS-val SPED Nordland
@@
Frozen 0 8.35 90.8 96.1 90.1 95.6 74.3 86.3
PartialTuned-2 14.18 15.54 90.8 95.1 89.0 95.4 79.3 89.6
PartialTuned-4 28.36 23.63 91.1 95.4 86.0 92.9 74.1 85.5
FullTuned 86.58 54.40 88.0 95.4 83.0 91.6 72.3 84.7
Adapter [25] 14.18 40.07 91.1 96.2 89.6 94.6 88.8 95.6
MCAdapter [24] 9.16 39.30 91.8 96.4 90.9 95.1 86.5 93.8
LoPA (ours) 0.08 10.09 92.0 96.6 92.4 95.9 88.3 95.3
Effect of LoPA. Next, to verify the effectiveness and efficiency of the proposed LoPA method, we compare it with other popular fine-tuning/adaptation methods, including PartialTuned (only fine-tuning the last few blocks of the backbone while keeping the others frozen), FullTuned (training the entire backbone), Adapter (as implemented in SelaVPR [25]) and MultiConvAdapter (as implemented in CricaVPR [24]). We use DINOv2-base as the backbone and set the frozen backbone without any adjustment as baseline. Besides, we consistently use our EDTformer as the aggregator. The results are presented in Table VI. In terms of efficiency, LoPA is an extremely lightweight network, solely introducing 0.08M trainable parameters, fewer than 1% of the parameters compared to other methods. In addition, compared to the frozen baseline, it only adds 1.74GB training memory, 10 times less than the PETL methods used in SelaVPR [25] and CricaVPR [24]. Amazingly, it saves over 40GB training memory usage compared to FullTuned. This demonstrates that LoPA is both parameter-efficient and memory-efficient, which makes it highly appropriate for VPR scenarios with constrained computational resources. In terms of accuracy, LoPA also significantly improves the performance of DINOv2 in VPR, while other methods result in minimal improvements or even a decline. Specifically, without any adjustment to DINOv2, our EDTformer remains competitive against some SOTA methods, such as MixVPR [33] and EigenPlaces [45]. Building on this, the LoPA further improves the baseline with 1.2%, 14.0% and 2.3% absolute R@1 on MSLS-val, Nordland and SPED. In contrary, PartialTuned yields minimal performance improvement on MSLS-val and Nordland. FullTuned, on the other hand, results in an overall performance decline. The PETL methods, i.e., Adapter and MultiConvAdapter, improve the recognition performance of the model on the MSLS-val and Nordland datasets. Yet these methods all lead to a decrease in R@5 on SPED. Comprehensively, the features output by DINOv2 with LoPA are more suitable for the sequent EDTformer to form the global representation, thereby achieving better performance at a minimal cost.
@@
TABLE IX: Ablation on FFN in the decoder. DINOv2-base with LoPA (rank set to 4) is used for feature extraction. The “w FFN” indicates that FFN is included, while “wo FFN” means FFN is removed.
Method Params MSLS-val SPED Pitts30k Nordland
@@
w FFN 15.76 91.9 96.6 90.8 96.0 92.8 96.6 84.2 92.7
w/o FFN 9.46 92.0 96.6 92.4 95.9 93.4 97.0 88.3 95.3
TABLE X: Ablation on the rank of LoPA. We extensively explore how different ranks impact the performance. DINOv2-base is used as the backbone with two stacked decoder blocks. “w/o LoPA” means only using frozen DINOv2 without any fine-tuning. The overall performance is best with rank set to 4.
@@
Figure 6: Visualization of learned queries. The left images with a green border show the original input images and their right counterparts with an orange border represent the attention weights of each learned query. It is worth noting that different queries prioritize different areas in the images, but they all focus on the temporarily constant and discriminative regions. On the contrary, for variable elements, these queries tend to discard them. For instance, in the first, second and last examples, all queries pay attention to the discriminative and invariant buildings and trees, while dynamic pedestrians and vehicles in the middle of the road are ignored. Thus, our method can retrieve the correct place images even if there are various changes in conditions.
@@
Effect of the FFN in the decoder block. We simplify the transformer decoder block by removing the feedforward network (FFN) to build the EDTformer for higher efficiency. Here, we conduct a simple experiment to explain why we do that. We consistently utilize two decoder blocks to construct EDTformer with 64 learnable queries. The results presented in Table IX demonstrate that adding FFN does not enhance the overall recognition performance. Instead, it leads to an obvious performance decline on multiple benchmark datasets, such as a 4.1% R@1 decline on Nordland. We think that this is because the large number of trainable parameters introduced by the FFN hinders the model from learning and aggregating effective features. Meanwhile, it also increases the complexity of the model and introduces additional memory overhead. Therefore, using a simplified decoder can improve efficiency and get better results, which is advantageous without any drawbacks.

### v2

@@ III-C Low-rank Parallel Adaptation
The foundation model DINOv2 recently has become a popular backbone for feature extraction in VPR, generally accompanied by a fine-tuning/adaptation for better performance. CricaVPR [3] and SelaVPR [4] have used the PETL methods (i.e., inserting tunable adapters into the frozen encoder blocks) to adapt DINOv2, as shown in Fig. 3 (b). However, they are efficient only in terms of parameters, not in training time and memory usage. Here, we conduct a simple theoretical derivation to explain this point. For a network with $L$ blocks, each having output ${z}_{i}$, a large number of frozen parameters $\theta_{i}$ and a few trainable parameters $\psi_{i}$, our goal is to minimize a loss function $\mathcal{L}$ based on stochastic gradient descent. Concretely, to update $\psi_{i}$ during backpropagation, we first need to compute $\frac{\partial\mathcal{L}}{\partial\psi_{i}}=\frac{\partial\mathcal{L}}{\partial z_{i}}\frac{\partial z_{i}}{\partial\psi_{i}}$. However, $\frac{\partial\mathcal{L}}{\partial z_{i}}$ cannot be computed directly, because $\mathcal{L}$ is usually calculated using the final output of the network (i.e, $z_{L}$). Therefore, it is necessary to resort to the chain rule as follows, which is the core of the famous backpropagation algorithm
$\frac{\partial\mathcal{L}}{\partial\psi_{i}}=\frac{\partial\mathcal{L}}{\partial z_{i}}\frac{\partial z_{i}}{\partial\psi_{i}}=\underbrace{\frac{\partial\mathcal{L}}{\partial z_{L}}\frac{\partial z_{L}}{\partial z_{L-1}}\cdots\frac{\partial z_{i+1}}{\partial z_{i}}}_{\text{intermediate activations}}\frac{\partial z_{i}}{\partial\psi_{i}}.$ (10)
Although most parameters (i.e., $\theta_{i}$) in each block are frozen, it is inevitable to compute gradients about the intermediate activations to assist in updating these few trainable parameters (i.e., $\psi_{i}$), as shown in Eq. (10), which incurs huge memory overhead and additional running time.
To tackle the above issue, we propose the Low-rank Parallel Adaptation method inspired by previous studies [37, 38, 39]. Instead of inserting tunable parameters inside the backbone, we design a parallel network that does not require backpropagation gradients through the frozen backbone during training as shown in Fig. 3 (c). It is a lightweight and separate network, which directly takes the intermediate features produced by each encoder block of the backbone as inputs and refines them progressively to output more accurate representations.
@@
Notably, our LoPA adapts the foundation model for the VPR task in a parameter- and memory-efficient way, thus providing more powerful features for subsequent aggregator and enhancing the robustness of the entire model. In other words, it is also applicable to other VPR methods when using the foundation model, instead of being tailored only for our EDTformer.
@@ III-D Training Strategy
The proposed model is trained on the GSV-Cities [66] dataset following its standard framework. GSV-Cities contains 560k images with highly accurate labels depicting 67k different places. Specifically, each place includes a set of images that describe the same location with different viewpoints and conditions. Besides, a unique label is assigned to each place. That is, the place images in GSV-Cities are divided into a restricted number of categories. Based on this characteristic, we apply the multi-similarity (MS) loss [54] function with an online hard mining strategy. The MS loss is formulated as
@@ IV-A Datasets and Evaluation Metrics
To demonstrate the effectiveness of our method, we conduct the experiments on multiple benchmark datasets, which present a diverse set of challenges encountered in the real world. Table I provides a concise summary of these datasets. Pitts30k [67], extracted from Google Street View, shows significant changes in viewpoint. MSLS [6] is collected from 30 major cities across six continents over nine years period, covering all seasons and encompassing diverse weather, cameras, daylight conditions, and structural settings. Tokyo24/7 [7] presents severe lighting (day/night) variations. SPED [8] is collected from surveillance cameras and consists of various low-quality images. Nordland [68] is captured using a front-mounted train camera, spanning all four seasons. AmsterTime [9] exhibits substantial image modality variations, using historical grayscale images as queries and contemporary color images as database. SVOX [10] is a cross-domain VPR dataset gathered under diverse weather and lighting conditions. We primarily utilize the three most challenging query subsets: SVOX Night, SVOX Rain and SVOX Sun. All these datasets are organized following the widely-used visual geo-localization benchmark [11].
@@ IV-B Implementation Details
TABLE III: Comparison to SOTA Methods on More Challenging Datasets. The Best Is Highlighted in Bold and the Second Is Underlined.
@@ IV-C Comparison with State-of-the-Art Methods
In this section, we compare our proposed method with a wide range of existing SOTA VPR algorithms, including ten single-stage methods using global features for direct retrieval: NetVLAD [19], SFRS [49], CosPlace [23], MixVPR [1], EigenPlaces [2], AnyLoc [33], CricaVPR [3], SuperVLAD [69], BoQ [5] and SALAD [24], as well as three superior two-stage methods: Patch-NetVLAD [21], $R^{2}$ormer [56] and SelaVPR [4]. Notably, our work uses the same training dataset as SALAD and BoQ, i.e., GSV-Cities. Additionally, the latest works CricaVPR, SuperVLAD, SALAD, BoQ and SelaVPR all leverage the foundation model DINOv2 as the backbone (SelaVPR using DINOv2-large, while others using DINOv2-base) and achieve the SOTA performance on multiple benchmarks. Thus, we also apply the DINOv2-base. Table II shows the quantitative results on MSLS, Tokyo24/7, SPED and Pitts30k. Our method achieves the best R@1/R@5/R@10 performance on almost all datasets.
@@
Results analysis. SALAD, BoQ, SelaVPR and our method all achieve superior performance on these datasets. Especially on Tokyo24/7, which shows severe illumination changes, SALAD, BoQ and SelaVPR achieve 94.6% R@1, 95.2% R@1 and 94.0% R@1 respectively. However, our method consistently improves performance on Tokyo24/7, achieving an incredible 97.1% R@1. This improvement stems from the robust and discriminative global descriptors produced by our model. The MSLS dataset is highly challenging, covering different seasons, weather and illumination conditions, various camera types and viewpoints, as well as different levels of dynamic objects presented in the scenes. Nevertheless, our method achieves 78.4% R@1, 89.8% R@5 and 91.9% R@10 on MSLS-challenge, showing significant advantages and outperforming other global-retrieval-base methods and two-stage methods by a considerable margin. Although our method only achieves 92.0% R@1 on MSLS-val, 0.2% lower than SALAD, the dimensionality of our global descriptors is less than half that of SALAD. This indicates that our global descriptors require less storage space and can achieve higher retrieval efficiency. In addition, we also get an overall better performance on the SPED dataset, demonstrating the high robustness of our method to process low-quality place images. On the Pitts30k dataset, which shows severe viewpoint changes, our EDTformer achieves a competitive performance, surpassing BoQ and SALAD by 1.5% and 0.9% on R@1 respectively. Although our EDTformer is slightly behind CricaVPR and SuperVLAD on individual datasets, this is because CricaVPR and SuperVLAD both use the cross-image encoder to combine multiple query images from the same place in a batch to achieve better performance. As the number of query images in the batch (i.e., batch size) decreases, especially when it becomes 1, their performance drops significantly, as shown in Table II CricaVPR-1 and SuperVLAD-1.
@@
To further evaluate the generalization of our method in some extreme scenarios, we conduct extensive experiments on other challenging datasets: Nordland, AmsterTime, SVOX Night, SVOX Rain and SVOX Sun. The results shown in Table III demonstrate the powerful capability of our method to effectively tackle the VPR task in difficult scenes. On the Nordland dataset, which exhibits significant variations in seasons and illumination, our EDTformer gets the best performance, even outperforming the two-stage SelaVPR. Additionally, our method surpasses other methods on the AmsterTime dataset by a large margin. Specifically, it outperforms the second by 10.9%, 8.2% and 6.7% on R@1/R@5/R@10 respectively. This implies the high robustness of our method to handle image modality variations in the datasets both containing grayscale and colorful images. On the SVOX Night, SVOX Rain and SVOX Sun datasets, our EDTformer achieves a relatively perfect performance (e.g., 98.7% R@1, 99.8% R@5 and 99.8% R@10 on SVOX Rain), which shows that it can be well applied to these special scenarios. In summary, the results on these challenging datasets further highlight the robust generalization capability of our model.
@@
Figure 4: The R@1 and inference time comparison of different single-stage methods on Pitts30k. We consistently measure the inference time on an NVIDIA GeForce RTX 4090 GPU.
@@
TABLE IV: The Comparison of SALAD, BoQ, SuperVLAD and EDTformer in Training Memory Usage and Training Time. We Train the Four Models on GSV-Cities with Batch Size Set to 36 and Measure the Metrics on an NVIDIA GeForce RTX 4090 GPU.
Metric SALAD BoQ SuperVLAD EDTformer
Training Memory Usage (GB) 14.81 15.21 10.23 5.72
Training Time (min / epoch) 11.21 10.67 10.57 8.77
Moreover, our method also exhibits an advantage in terms of efficiency in multiple aspects. Fig. 4 exhibits the R@1 performance on the Pitts30k dataset and the inference time of a single image about seven global-retrieval-based methods, including SFRS, MixVPR, SALAD, BoQ, SuperVLAD, SelaVPR (global) and our EDTformer. MixVPR uses the CNN model (ResNet-50) as the backbone and proposes a feature mixing method to get global descriptors, which achieves the fastest inference speed. Although SFRS also leverages the CNN model (VGG-16), it applies PCA to reduce the dimensionality of the global descriptors, which is time-consuming. The other methods are all based on the vision foundation models. Specifically, SelaVPR (global) adopts DINOv2-large as the backbone, resulting in the slowest inference speed (16.95 ms). Among SALAD, BoQ, SuperVLAD and our EDTformer, all of which use DINOv2-base as the backbone, our method achieves the fastest inference speed (9.16 ms). In addition, we also provide the training memory consumption and training time per epoch for these four methods, as shown in Table. IV. Our training speed is approximately 15% faster. Besides, our training memory usage is only about half that of SuperVLAD and one-third that of SALAD and BoQ. In summary, our method is efficient both in training and inference while achieving a superior recognition performance.
@@ IV-D Ablation Studies
In this section, we first conduct a series of ablation experiments to demonstrate the effectiveness of our proposed EDTformer and LoPA. Subsequently, we further explore the impact of some design details. Notably, we consistently use GSV-Cities with the MS loss for training in the ablation experiments. Unless otherwise specified, the dimensionality of the global features produced by our method is 4096.
@@
Effect of EDTformer. To validate the effectiveness of the proposed EDTformer, we compare it with current common aggregation methods, including GeM [20], NetVLAD [19], Conv-AP [66], MixVPR [1] and BoQ [5] (related to our method). The “GeM-linear” represents the use of the GeM aggregator, followed by a linear layer to adjust the dimensionality of the descriptors. Likewise, “NetVLAD-PCA” implies that a PCA layer is integrated after the NetVLAD aggregator to perform dimensionality reduction. To ensure the fairness of the experiments, we consistently use the frozen DINOv2-base as the backbone with LoPA (rank set to 4) for feature extraction. The results are presented in Table V. When the dimensionality of global descriptors is the same (4096-dim), our EDTformer achieves highly competitive performance and outperforms other common aggregators (e.g., GeM, NetVLAD, Conv-AP and MixVPR) by a large margin. Although BoQ achieves outstanding results, our method continues to improve the performance, especially on the Nordland dataset (obtaining an improvement of 4.5% in R@1). This is due to the fact that our EDTformer uses a simpler decoder structure to achieve more consistent feature aggregation. Notably, the dimensionality of the global descriptors (12288-dim) produced by BoQ is more than three times that of our global representations (4096-dim), which indicates our EDTformer can achieve higher retrieval efficiency than BoQ. Moreover, we simultaneously present the recognition performance of our global descriptors with varying dimensions (achieved by only adjusting the last linear layer in our pipeline). As the dimensionality of the descriptors decreases, the overall performance gradually declines. However, our 512-dim descriptors can still achieve an impressive 90.1% R@1 on MSLS-val, 89.3% R@1 on SPED, 91.7% R@1 on Pitts30k and 71.1% R@1 on Nordland, surpassing the performance of the 768-dim descriptors output by GeM and remaining competitive against some SOTA methods such as SFRS [49], CosPlace [23] and EigenPlace [2]. Therefore, our EDTformer is also well-suited for some VPR scenarios that urgently require low-dimensional descriptors.
TABLE VI: Ablation on Different Fine-tuning/Adaptation Methods. “PartialTuning-2” and “PartialTuning-4” Respectively Denote Only Fine-tuning the Last 2 and 4 Encoder Blocks of DINOv2. “MCAdapter” Is the Abbreviation of MultiConvAdapter. The Trainable Parameters Introduced by Different Methods Are Reported. We Use an NVIDIA A800 GPU to Measure the Training Memory Usage with Batch Size Set to 72.
Method Params Memory MSLS-val SPED Nordland
@@
FrozenDINOv2 [33] 0 8.35 90.8 96.1 90.1 95.6 74.3 86.3
PartialTuning-2 [5] 14.18 15.54(+7.19) 90.8 95.1 89.0 95.4 79.3 89.6
PartialTuning-4 [24] 28.36 23.63(+15.28) 91.1 95.4 86.0 92.9 74.1 85.5
FullTuning 86.58 54.40(+46.05) 88.0 95.4 83.0 91.6 72.3 84.7
Adapter [4] 14.18 40.07(+31.72) 91.1 96.2 89.6 94.6 88.8 95.6
MCAdapter [3] 9.16 39.30(+30.95) 91.8 96.4 90.9 95.1 86.5 93.8
LoPA (ours) 0.08 10.09(+1.74) 92.0 96.6 92.4 95.9 88.3 95.3
Effect of LoPA. Next, to verify the efficiency and effectiveness of the proposed LoPA, we compare it with other common fine-tuning/PETL methods in VPR, including PartialTuning used in BoQ [24] and SALAD [24] (only fine-tuning the last few encoder blocks of the backbone), FullTuning (training the whole backbone), Adapter (as implemented in SelaVPR [4]) and MultiConvAdapter (as implemented in CricaVPR [3]). We use DINOv2-base as the backbone and set the frozen backbone without any adjustment as baseline. Besides, we consistently use our EDTformer as the aggregator. The results are presented in Table VI. (1) In terms of efficiency, LoPA is an extremely lightweight network, solely introducing 0.08M trainable parameters, fewer than 1% of the parameters compared to other methods. In addition, compared to the frozen baseline, it only results in 1.74GB additional training memory usage, approximately 1/4 of the PartialTuning-2 method and even 10 times less than the PETL methods used in SelaVPR and CricaVPR. Amazingly, it saves over 40GB training memory usage compared to FullTuning. This demonstrates that LoPA is parameter- and memory-efficient, which makes it highly appropriate for VPR scenarios with constrained computational resources. (2) In terms of accuracy, LoPA also significantly improves the performance of DINOv2 in VPR, while other methods result in minimal improvements or even a decline. Specifically, without any adjustment to DINOv2, our EDTformer remains competitive against some SOTA methods, such as MixVPR [1] and EigenPlaces [2]. Building on this, our LoPA further improves the baseline with 1.2%, 14.0% and 2.3% absolute R@1 on MSLS-val, Nordland and SPED. In contrary, PartialTuning yields minimal performance improvement on MSLS-val and Nordland. FullTuning, on the other hand, results in an overall performance decline. Although the PETL methods, i.e., Adapter and MultiConvAdapter, improve the recognition performance of the model on the MSLS-val and Nordland, they lead to a decrease in R@5 on SPED. Comprehensively, the features output by DINOv2 with LoPA are more suitable for the sequent EDTformer to form global representations, which is one of the reasons why our model performs better than other methods (e.g., BoQ, SALAD, etc).
@@
TABLE IX: Ablation on FFN in the Decoder Block. The Total Trainable Parameters, R@1 and R@5 Performance Are Reported.
FFN LoPA Params MSLS-val SPED Pitts30k Nordland
@@
$\times$ 16.59 90.5 96.2 89.6 95.1 91.6 96.4 74.9 86.9
$\times$ 10.29 90.8 96.1 90.1 95.6 91.8 96.3 74.3 86.3
16.68 91.9 96.6 90.8 96.0 92.8 96.6 84.2 92.7
$\times$ 10.38 92.0 96.6 92.4 95.9 93.4 97.0 88.3 95.3
TABLE X: Ablation on the Rank of LoPA. We Extensively Explore How Different Ranks Impact the Performance. DINOv2-base Is Used as the Backbone. “w/o LoPA” Means Only Using Frozen DINOv2 Without Any Fine-tuning. The Overall Performance Is Best with Rank Set to 4.
@@
TABLE XI: Ablation of LoPA on Other VPR Methods. We Consistently Use the Frozen DINOv2-base Backbone.
Method LoPA Pitts30k MSLS-val Nordland SPED
R@1 R@5 R@1 R@5 R@1 R@5 R@1 R@5
GeM $\times$ 78.4 89.2 36.4 47.0 12.1 21.3 58.5 75.3
88.5 94.4 85.7 94.6 32.3 47.1 76.1 88.5
NetVLAD $\times$ 82.7 92.8 55.3 67.7 31.1 46.2 80.2 92.3
92.0 96.3 89.7 95.5 71.7 84.8 88.8 95.1
SALAD $\times$ 91.5 96.1 88.9 95.4 62.0 75.5 88.3 94.4
92.5 96.4 91.6 96.2 84.1 92.5 91.3 96.2
@@
Figure 6: Visualization of learnable queries. The first column (a) with a green border shows the original input images, the middle 2-5 columns (b) with an orange border represent the attention weights of a single learnable query, and the last column (c) with a blue border denotes the union of all query attention weights. It is worth noting that different queries prioritize different areas in the images, but they all focus on the temporarily constant and discriminative regions. On the contrary, for variable elements, these queries tend to discard them. Concretely, in the first two examples, all queries pay attention to the discriminative and short-term unchanged buildings, trees, and special signs on the ground while dynamic pedestrians and vehicles are basically ignored. The third and fourth examples are taken from almost the same location but with different illumination and seasons. Nevertheless, our learnable queries consistently focus on the condition-invariant house, bridge, and special landmark. When the place images (i.e., the last two examples) come from different perspectives and span a very long time, our method can pay attention to the discriminative and unchanged regions that they commonly possess.
@@
Effect of the FFN in the decoder block. We simplify the standard transformer decoder block by removing the feedforward network (FFN) to build our EDTformer for higher efficiency. Here, we conduct the corresponding experiments to explain why we do that. We consistently utilize two decoder blocks to construct EDTformer with 64 learnable queries. For comprehensiveness, we conduct the experiments with and without LoPA. The results are presented in Table IX. Without LoPA to adapt the frozen DINOv2, our simplified decoder performs similarly to the standard transformer decoder. With our LoPA, however, using the simplified decoder can achieve a better overall recognition performance, such as an absolute 4.1% R@1 improvement on Nordland. This indicates the features produced by DINOv2 with LoPA are more suitable for the sequent EDTformer consisting of our simplified decoder blocks. Besides, our simplified decoder introduces less trainable parameters and can achieve higher efficiency compared to the standard transformer decoder. Therefore, using our simplified decoder block is a better choice, which is advantageous without any drawbacks.